OpenStudy (samanthagreer):

Which step should be completed first to solve the equation ? A. Add 14 to both sides. B. Divide both sides by 3. C. Divide both sides by −3. D. Subtract 9 from both sides.

OpenStudy (shinalcantara):

first thing to do is subtract 9 from both sides of the equation 9 - 9 - 3q = 14 - 9 -3q = 5 the main purpose for this is to have the constants lying on the same side of the equation, leaving the variable alone on the other side